OpenStudy (anonymous):

y"+αy'+βy = 0. let sum of the roots of its auxiliary eq be 94 , product of the roots is 2013. how to solve the DE.

OpenStudy (amistre64):

should we assume an e^rx solution, or some x^r solution?

OpenStudy (amistre64):

the key might be in getting a system of equations from roots

OpenStudy (amistre64):

r^2 + ar + b = 0 when -a +- sqrt(a^2 -4b) = 0

OpenStudy (anonymous):

should we use the sum n the product of the roots ?

OpenStudy (amistre64):

yes, but im not too sure if we need to run thru the auxilary stuff r1 + r2 = 94 r1 * r2 = 2013 should develop roots that fit the bill

OpenStudy (dumbcow):

based on sum = 94 and product = 2013 we can conclude that a = 94 b = 2013

OpenStudy (amistre64):

(47-n) +(47+n) = 94 47^2 - n^2 = 2013 solve for n :)

OpenStudy (amistre64):

it looks like n = 14 47+14 = 61 47-14 = 33

OpenStudy (amistre64):

i wonder if these are suitable roots ....

OpenStudy (anonymous):

okay.. but if i use the info from the question given, \[\alpha + \beta = 94 , \alpha \beta = 2013\] how to get the value of alpha and beta ?

OpenStudy (amistre64):

alpha and beta would be a part of the auxil eq such that\[r=\frac{-a\pm\sqrt{a^2-4b}}{2}=61,33\]

OpenStudy (amistre64):

-a/2 = middle of 61 and 33; 47 a = -94

OpenStudy (amistre64):

\[\frac{\sqrt{94^2-4b}}{2}=14\] \[\sqrt{94^2-4b}=28\] \[94^2-4b=28^2\] \[\frac{94^2-28^2}{4}=b=2013\] just like dumbcow prophecised :0

OpenStudy (anonymous):

I've seen this exact question here before: http://openstudy.com/study#/updates/51a0f70ae4b0e9c70c332ec0 @amistre64 and @dumbcow have got it covered, though.

OpenStudy (anonymous):

okay..thanks for your help @amistre64 @dumbcow @SithsAndGiggles ;)